The price of a stock is $75 per share. The stock is not expected to pay any cash dividends over the

next year. The returns on the stock have a normal probability distribution with expected value of 15%

and a standard deviation of 40%.

(a) What is the probability that the price of the stock will be $25 or less one year from now?

(b) What is the probability that the price of the stock will be between $75 and $100 one year from

now?

(c) What is the probability that the price of the stock will be exactly $25 one year from now?

(d) What is the minimum amount you can lose on the stock on a per share basis with the probability

of 10%?
Answer #1

1.
=NORMDIST(25/75-1,15%,40%,TRUE)=2.06%

2.
=NORMDIST(100/75-1,15%,40%,TRUE)-NORMDIST(75/75-1,15%,40%,TRUE)=32.28%

3.
0

4.
=NORMINV(0.1,15%,40%)*75=27.1965